OpenStudy (anonymous):

To use mental math and compare 5/8 to 1/2, which fraction would you use?

OpenStudy (anonymous):

I can't seem to know your thought. lol. I would use 1/2

OpenStudy (anonymous):

denominator is 8\[\frac{1}{2}=\frac{1*4}{2*4}=\frac{4}{8}\]\[\frac{5}{8}>\frac{4}{8}\]than 5/8>1/2

OpenStudy (anonymous):

Yes, not sure what you mean, but when adding fractions you can switch it so they both have same denominator, makes it easier to add. it becomes (5/8)+(4/8)

OpenStudy (anonymous):

Nancy is fast.

OpenStudy (anonymous):

ty